### v1

# Sleeping Kelly is a Thirder

Abstract
The Sleeping Beauty problem was presented by Elga to highlight the role of probabilities in situations with imperfect recall. One approach to solving the Sleeping Beauty problem is to allow Sleeping Beauty to make decisions based on her beliefs, and then characterize what it takes for her decisions to be “rational”. In particular, she can be allowed to make monetary bets based on her beliefs, with the assumption that she wants to gain wealth rather than lose it. However, this approach is often coupled with the assumption that Sleeping Beauty should maximize the expected value of her bets. Here, I argue instead that it is rational for Sleeping Beauty to maximize the growth rate of her wealth using the Kelly Criterion, which leads us to the “thirder” position. Furthermore, this position is shown to be “rational” by Dutch book arguments. If Sleeping Kelly only accepts bets that have a growth rate greater than 1 as a “thirder” then she is not vulnerable to Dutch books. By contrast, if Sleeping Beauty takes the “halfer” position, she is vulnerable to Dutch books.
1 Problem Statement
The Sleeping Beauty problem was presented as follows by Elga (2000) [^1] A version of the problem is thought to be first proposed by Zuboff (1990):
“Some researchers are going to put you to sleep. During the two days that your sleep will last, they will briefly wake you up either once or twice, depending on the toss of a fair coin (Heads: once; Tails: twice). After each waking, they will put you to back to sleep with a drug that makes you forget that waking. When you are first awakened, to what degree ought you believe that the outcome of the coin toss is Heads?”
SleepFlipWakeWakeWakeSundayMondayTuesdayHT
2 Background
The two most common answers are to assign probabilities of $\frac{1}{2}$ (“halfer” position) and $\frac{1}{3}$ (“thirder” position) to the coin having landed heads. Elga (2000) argued for the thirder position, while Lewis (2001) prominently argued for the halfer position, but debate continues. According to Bostrom (2007), “neither side seems to have gained a decisive advantage.”
Upon waking Sleeping Beauty, if we ask her to assign a probability to the coin landing heads this leaves open the question of her objective. In other words, how bad is it to say $\frac{1}{3}$ if the “true” answer is $\frac{1}{2}$ and vice versa? Probabilities do not have inherent physical meaning, but objectives can. Allowing Sleeping Beauty to make decisions forces her beliefs to make contact with reality, and we can create criteria to evaluate these decisions. Here enters decision theory.
The simplest type of decision to evaluate is perhaps a monetary bet, in which the basic assumption that Sleeping Beauty would prefer to increase her wealth. The trouble with using monetary bets is that one must make assumptions about Sleeping Beauty’s betting strategy based on the probabilities she assigns to different events. The existing literature from decision theory on the Sleeping Beauty problem uniformly makes the same assumption: that it is rational for Sleeping Beauty to maximize expected value.
“It is a platitude among decision theorists that agents should choose their actions so as to maximize expected value. But exactly how to define expected value is contentious.”- Briggs (2010)
Arguably the most substantive analysis of the Sleeping Beauty problem since Elga (2000) is Halpern (2004), which focuses on synchrony and asynchrony. However, these arguments are complicated, and as we will see, as soon as Halpern (2004) considers betting arguments the reasoning falls flat, precisely because it focuses on expectation values of bet payoffs.
In many betting-based arguments, the expected value (ensemble average) of an observable is conflated with the growth rate. If Sleeping Beauty is offered a sequence of gambles, or expects to be offered a sequence of gambles, it is rational for her to maximize the (exponential) growth rate of her wealth, which is not the same as maximizing expected value (Kelly, 1956). But to optimize her growth rate, Sleeping Beauty needs to have some information about what fraction of her wealth she is wagering. [^2] No poker player would ever determine how much to bet on a hand without regard for how much money they have left.
The conflation of expectation value with the growth rate leads to erroneous conclusions, such as that if Sleeping Beauty is offered bets with 1:1 odds on a fair coin flip then, “she will be indifferent between accepting them and not accepting them” (Conitzer, 2015). With any finite amount of initial wealth and any fixed bet size this would cause Sleeping Beauty’s wealth to follow an unbiased random walk over repeated sequences awakenings or repeated experiments, which guarantees eventual bankruptcy (Pearson, 1905).
We now examine a version of the Sleeping Beauty problem with wagers that create multiplicative dynamics rather than additive dynamics. All of Sleeping Beauty’s wagers are expressed as a fraction of the money she has, rather than a fixed bet size. In this situation, Sleeping Beauty maximizes the growth rate of her wealth using the Kelly Criterion (Kelly, 1956).
3 Problem Setup: Sequential Betting
You are Sleeping Beauty. You begin the experiment with a certain amount of money which you can use to make wagers during the experiment. Some researchers are going to put you to sleep on Sunday. Before you are put to sleep you have the option to bet any fraction of your wealth $[0,1]$ on the outcome of a fair coin flip at 1:1 odds. You do not have access to leverage, so you cannot bet more than 100% of your wealth. The coin will be flipped after you are put to sleep and you will not be shown the outcome until the experiment is over on Wednesday. If the coin lands heads you will be woken up on Monday. If the coin lands tails you will be woken up on Monday and Tuesday. Each time you are woken up you will have the option to wager any faction of your money on the outcome of the coin flip at 1:1 odds before being put back to sleep. You will not be informed of how much money you have at any point, so all of your wagers must be given as a fraction of your wealth rather than a monetary value. Each time you are put back to sleep your memory will be erased so you have no recollection of being woken or your previous wager if there was one. Your goal is to determine what fraction of your wealth to wager on the coin landing heads or tails before being put to sleep and each time you are woken.
4 Expected Value versus Growth Rate
What is the optimal betting strategy in the Sleeping Beauty problem where bets are a fraction of your wealth? Let $1\leq a\leq 1$ be the fraction of wealth you wager on the coin landing tails before being put to sleep, and let $1\leq b\leq 1$ be the fraction of wealth you wager on the coin landing tails each time you are woken during a run of the experiment. If $a<0$ or $b<0$, this means wagering on the coin landing heads instead of tails.
4.1 Expected Value
If your goal were to maximize the expected value of your wealth at the end of each experiment, then you would choose $a$ and $b$ to maximize the value $\frac{(1-a)(1-b)+(1+a)(1+b)(1+b)}{2}$. This function is maximized by setting $a=b=1$; going all-in on every wager, to get an expected value of 4 times your original wealth [^3] This is a polynomial in two variables, so the maxima are found by taking the partial derivatives to get the critical points, and second derivatives to determine the maxima.. After the experiment is over, you will either have no money left or 8 times your original endowment.
The idea of betting 100% of your wealth on a coin flip, before being put to sleep or when woken up, should feel wrong. It is. If you run the experiment many times in sequence as a single individual, you will eventually go bankrupt with probability 1. Eventually you will lose a bet, and from then on there is no money left to wager.
4.2 Growth Rate and the Kelly Criterion
Now suppose we are going to repeat the experiment many times in sequence, what is the optimal betting strategy?
Theorem 1.
Sleeping Beauty achieves the maximum growth rate for her wealth as a “thirder” who sizes her bets using the Kelly Criterion.
Suppose we run the experiment twice in a row, and the coin lands heads once and tails once. Then your wealth at the end of the two experiments is (1a)(1b)(1+a)(1+b)(1+b)(1-a)(1-b)(1+a)$(1+b)$ times your original wealth.33footnotemark: 3 To maximize this value we must choose $a=0$ and $b=\frac{1}{3}$.
Of course, if we run the experiment only twice we are not guaranteed to get one heads and one tails from the coin flips. However, as the number of experiments tends to infinitely, we expect half of the flips to be heads and half to be tails across the experiments, by the law of large numbers. The wealth multiplier is therefore $\sqrt{(1-a)(1-b)(1+a)(1+b)(1+b)}$ per experiment. Thus, $a=0$, $b=\frac{1}{3}$ is the optimal strategy for long-term wealth maximization. You do not bet any amount before being put to sleep, but each time you are woken it is best to wager $\frac{1}{3}$ of your wealth that the coin landed tails.
Figure 1: Wealth multiplier for wager fractions $a$ and $b$.
This answer at least passes the smell test. One should not bet any fraction of your wealth on a fair coin toss before the experiment begins. Imagine choosing $a>0$ and $b=0$, so you are simply successively betting on successive coin flips. Then by repeating the experiment many times in sequence, your wealth will eventually converge to 0 for any value of $0<a\leq 1$, because (1a)(1+a)<1(1-a)(1+a)<1.
From here on, we will use the wealth multiplier instead of the actual growth rate because it is simpler to work with. We refer to Peters and Adamou (2025) for definitions of wealth multipliers and growth rates.
Inferring Probability from the Kelly Criterion
We can now infer the probability implicitly assigned to the coin landing on heads using the Kelly Criterion, which maximizes the growth rate, or equivalently the multiplier, of wealth under multiplicative dynamics (Thorp, 1975; Kelly, 1956; Peters and Adamou, 2025). Given 1:1 odds on any bet when woken, the Kelly Criterion says to bet a fraction of your wealth b=p(1p)b=p-(1-p), where $p$ is the probability when woken that the coin landed on tails. By setting $b=\frac{1}{3}$, we infer $p=\frac{2}{3}$. In other words, you must implicitly assign a probability $\frac{2}{3}$ to the coin landing tails, and $\frac{1}{3}$ to the coin landing heads, each time you are woken. Similarly, since $a=0$ we infer $p=\frac{1}{2}$ before being put to sleep. Sleeping Kelly is a “thirder.”
5 Dutch Books
We saw in Section 4.2 that the “thirder” position allows Sleeping Beauty to maximize the growth rate of her wealth when betting according to the Kelly Criterion. In that setup, Sleeping Beauty got to decide what fraction of her wealth to wager, and whether to wager it on heads or tails. We now look at a setting in which Sleeping Beauty’s strategy space is more restricted. Sleeping Beauty will be given an offer to bet on the outcome of the coin flip before she is put to sleep and each time she is woken, but the experimenter will determine the wager, and Sleeping Beauty only has the option to accept or reject the bet. When put back to sleep, she will have no memory of what previous bets she has been offered or whether she accepted them. We require that the same bet be offered whenever she is woken, and that it be determined before the experiment begins, so the offered bet cannot depend on the coin’s outcome or what day it is.
The question we want to answer is whether Sleeping Beauty can be subjected to a Dutch Book. As Conitzer (2015) explains,
“One decision-theoretic approach to answering this question is to attempt to construct a Dutch book to which one of the two positions is vulnerable. A Dutch book is a set of bets that the agent in question would all accept individually, but that together ensure that the agent incurs a strict loss overall. In a diachronic Dutch book, the bets are offered at different times[.]”
5.1 Dutch Book of Expected Value Maximizer
Suppose that Sleeping Beauty were to accept any bet with a positive expected value. Different Dutch books have been constructed for Sleeping Beauties who are “evidential decision theorists” and “causal decision theorists” that maximize expected value. Consider the following basic example against the halfer position from Hitchcock (2004):
Before going to sleep, Sleeping Beauty is offered a wager at almost 1:1 odds that wins $15+\epsilon$ if the coin lands tails and loses 15-15 if the coin lands heads. Each time she is woken, she is offered a bet that wins $10+\epsilon$ if the coin landed heads, and loses 10-10 if the coin landed tails. Both bets have an expected value slightly greater than zero, so Sleeping Beauty is presumed to accept all bets she is offered. However, if the coin lands heads she ends up losing 5-5 and if the coin lands tails she also ends up losing 5-5 by the end of the experiment.
The key flaw here is that a rational sequential betting strategy for Sleeping Beauty cannot be defined without any information about how the size of the bet relates to the amount of money she has available. We want to see whether any Dutch book can be created for a version of Sleeping Beauty who is offered sequential wagers expressed as a fraction of her wealth, and who only accepts bets that increase the growth rate of her wealth, i.e., have a wealth multiplier greater than 1.
5.2 Invulnerability of Sleeping Kelly
We want to know whether Sleeping Beauty can be subjected to a Dutch book if she is only willing to accept bets that individually have a growth rate greater than 1, or equivalently with a wealth multiplier greater than 1; i.e., greater than just sitting on the cash and not betting it.
Suppose there is an event with $k$ possible outcomes $\{A_{1},\ldots,A_{k}\}$ with respective probabilities $p_{1},\ldots,p_{k}$. You are offered a bet where your wealth is multiplied by $(1+\alpha_{i})$ if event $A_{i}$ occurs for each $i\leq k$. If you require your wealth multiplier to be greater than one then this means you require $\prod\limits_{i\leq k}(1+\alpha_{i})^{p_{i}}>1$.
Theorem 2.
Sleeping Kelly cannot be subjected to a Dutch book if she is a “thirder” and only accepts bets with a wealth multiplier greater than 1.
Proof.
Let’s define the following variables which all lie in the range [1,1][-1,1]:
  • $\alpha_{OH}$ = fraction of wealth won from bet before going to sleep if coin lands heads
  • $\alpha_{OT}$ = fraction of wealth won from bet before going to sleep if coin lands tails
  • $\alpha_{WH}$ = fraction of wealth won from bet when woken up if coin lands heads
  • $\alpha_{WT}$ = fraction of wealth won from bet when woken up if coin lands tails
We assume that Sleeping Beauty will not accept a bet unless it has a wealth multiplier greater than one based on the probabilities she assigns. As a “thirder” Sleeping Beauty assigns a probability of $\frac{1}{2}$ to the coin landing tails before being put to sleep and probability $\frac{2}{3}$ to the probability of tails upon being woken.
We now have two conditions for whether Sleeping Beauty will accept a bet offered to her. Before sleeping she requires $(1+\alpha_{OH})^{1/2}(1+\alpha_{OT})^{1/2}>1$. Upon waking she requires $(1+\alpha_{WH})^{1/3}(1+\alpha_{OT})^{2/3}>1$ to accept the bet. We can simplify the exponents to get our two conditions for bet acceptance.
Bet Acceptance Conditions:
  • Before Sleeping: $(1+\alpha_{OH})(1+\alpha_{OT})>1$
  • Upon Waking: $(1+\alpha_{WH})(1+\alpha_{WT})^{2}>1$
To construct a Dutch book, in which Sleeping Beauty loses money regardless of the outcome of the coin flip, we would need a set of values such that the following two conditions hold.
Dutch Book Conditions:
  • Losing When Heads: $(1+\alpha_{OH})(1+\alpha_{WH})<1$
  • Losing When Tails: $(1+\alpha_{OT})(1+\alpha_{WT})^{2}<1$
If both of the Dutch book conditions hold, then by multiplying them together, we must also have the condition that the wealth multiplier is less than 1:
$(1+\alpha_{OH})(1+\alpha_{WH})(1+\alpha_{OT})(1+\alpha_{WT})^{2}<1.$
However, this is not possible. By reordering the terms, we see
$\underbrace{(1+\alpha_{OH})(1+\alpha_{OT})}_{>1}\underbrace{(1+\alpha_{WH})(1+\alpha_{WT})^{2}}_{>1}>1$
according to our bet acceptance conditions.
Thus, if Sleeping Beauty only accepts bets that individually have a wealth multiplier greater than 1, her overall growth rate must be greater than 1 and she cannot be subjected to a Dutch book. Her wealth can only grow or stay constant (if she rejects the offers). ∎
6 Vulnerability of Halfers
Suppose Sleeping Beauty only accepts bets with wealth multiplier greater than 1 but she assigns a probability $p=\frac{1}{2}$ to the coin landing heads each time she is woken. We show that these “halfers” are vulnerable to Dutch books.
Theorem 3.
Sleeping Beauty is vulnerable to Dutch books if she assigns a probability of $\frac{1}{2}$ to the coin landing heads or tails each time she is woken, even if she only accepts bets with a growth rate greater than 1.
Now we have a different set of bet acceptance conditions, but the same Dutch book conditions as before.
Bet Acceptance Conditions:
  • Before Sleeping: $(1+\alpha_{OH})(1+\alpha_{OT})>1$
  • Upon Waking: $(1+\alpha_{WH})(1+\alpha_{WT})>1$
Dutch Book Conditions:
  • Losing When Heads: $(1+\alpha_{OH})(1+\alpha_{WH})<1$
  • Losing When Tails: $(1+\alpha_{OT})(1+\alpha_{WT})^{2}<1$
Now a Dutch book is possible. For example we can offer the “halfer” the following bets:
  • $\alpha_{WH}=\frac{1}{2}$
  • $\alpha_{WT}=-\frac{1}{3}+\epsilon$
  • $\alpha_{OT}=1$
  • $\alpha_{OH}=-\frac{1}{3}-\epsilon$
Before sleeping, her wealth multiplier would be $\frac{4}{3}-\epsilon$ and upon waking her wealth multiplier would be $1+\epsilon$. However, if the coin lands heads then her wealth shrinks by a factor of $1-\epsilon$, and if the coin lands tails then her wealth shrinks by a factor of $\frac{8}{9}-\epsilon$.
7 Conclusion: Sleeping Kelly is a “thirder”
We approach the Sleeping Beauty problem by allowing Sleeping Beauty to place bets based on her beliefs. However, in contrast to canonical decision theory, we do not assume that Sleeping Beauty’s betting strategy is based on maximizing expected value. Instead, we assume that Sleeping Beauty aims to maximize the growth rate of her wealth, invoking the Kelly Criterion under multiplicative dynamics. This leads us to the conclusion that before the coin is flipped, Sleeping Beauty should bet based on a probability of $\frac{1}{2}$ that the coin lands tails, but each time Sleeping Beauty is woken she should bet based on a probability of $\frac{1}{3}$ that the coin lands tails.
Another approach to addressing the Sleeping Beauty problem is by examining whether Sleeping Beauty’s bets can be subjected to a Dutch book as a condition for rationality. We show that if Sleeping Kelly is a “thirder” and only accepts bets with a wealth multiplier greater than 1, she is not vulnerable to a Dutch book. However, “halfers” are vulnerable to Dutch books, even if they only accept bets they expect to have a wealth multiplier greater than 1.
Myopically maximizing the expected value of individual gambles is suboptimal for Sleeping Beauty, vulnerable to Dutch books, and arguably irrational.